OpenStudy (anonymous):

Can you solve the following inequality? one over x - 8 is less than or equal to 1 over x - 7 1 / x−8 ≤ 1 / x−7

OpenStudy (anonymous):

=>1/(x-8) - 1/(x-7)<=0 =>(x-7-x+8)/{(x-8)(x-7)} <= 0 => 1/(x-8)(x-7) <=0 this is only going to happen if (x-8)(x-7)<0 i.e 7<x<8